OpenStudy (anonymous):

You research the cost of a gallon of gasoline over several years to look for a trend. The table shows your data. The line of best fit uses years since 1980 as the input value. What is a line of best fit for the data? Based upon these prices, how much would you expect to pay in 2030?

OpenStudy (anonymous):

Year 1980 1984 1988 1992 1996 2000 2004 2008 2012 Price per Gallon $1.25 $1.27 $1.05 $1.28 $1.43 $1.68 $2.11 $3.72 $3.99

OpenStudy (anonymous):

y = 0.087x + 2.26; $6.44 y = 0.087x + 0.587; $4.76 y = 0.134x + 1.25; $7.68 y = 0.0347x + 0.856; $2.52

OpenStudy (amistre64):

what do you have as far as tools that you are allowed to use to determine a solution with?

OpenStudy (amistre64):

if it has to be done by hand, it can get messy if you can use a ti83 or similar, or an excel sheet, its becomes much simpler